# Dimension-reduced Reconstruction Map Learning for Parameter Estimation in Likelihood-Free Inference Problems

### v1

@@ 1 Introduction
Our approach shares a foundation with recent literature on what we shall call reconstruction map (RM) estimation. Rudi et al. (2022) consider parameter estimation for the FitzHugh–Nagumo model by learning the mapping from the sample space to parameter space using a deep NN trained on a large number of synthetic datasets generated from the model. Lenzi et al. (2023) further expand its application to intractable models with an example of parameter estimation for max-stable processes. Crucially, their approach suffers from the curse of dimensionality, i.e., its estimation performance degrades quickly as the data size grows. In an application for econometric models, Creel (2017) proposes to use informative statistics as input to train a neural network, the output of which can be used directly as an estimator, or as an input to subsequent classical or Bayesian inference estimation. Our work builds on this approach by establishing a systematic simulation-based Reconstruction Map-dimension Reduction (RM-DR) estimation method which overcomes the fundamental problem of degraded estimation performance with data dimension. We show that under certain assumptions, the resulting estimator is asymptotically equivalent to a Bayes estimator. Through multiple numerical experiments, we show that dimension reduction is essential for estimation from large datasets. We further propose a combined parameter estimation approach that utilizes the RM-DR as a starting point in a local optimization algorithm when the likelihood is available, providing an alternative to computationally costly global optimization methods.
@@ 2.1 Neural Network Models
The key building block of the NN is the neuron as shown in the left panel of Figure 1, which consists of $n^{in}$ input nodes/variables with the $i$th variable denoted as $x_{i}$ and associated weight parameter $w_{i}$, a bias parameter $b$, an activation function $f:\mathbb{R}\rightarrow\mathbb{R}$, and the neuron output $x^{out}$. Let $x^{in}$ and $w$ be vector representations of input variables and weights, respectively. The output $x^{out}$ is calculated as evaluating the activation function at the weighted sum of input variables,
@@
More complex models can be constructed using multi-layer NNs, where there is a sequence of hidden layers between input and output layer as shown in the right panel of Figure 1. For an L-layer NN, denote $A_{l}$ as the matrix of all parameters in the $l$th layer, and $f_{l}$ as the activation function in $l$th layer, for $l=1,\ldots,L$. Let $\omega:=(A_{1},\ldots,A_{L})$ be parameters of the NN represented as a sequence of matrices, and $\mathbf{N}(\cdot,\omega):\mathbb{R}^{n^{in}}\rightarrow\mathbb{R}^{n^{out}}$ be the vector-valued function representation of the NN,
@@ 2.2 Reconstruction Map (RM) Estimation
$\widehat{\theta}_{RM}(y)=\mathbf{N}(y,\omega^{*}),\;\text{ where }\omega^{*}=\underset{\omega\in\Omega}{\operatorname{argmin}}\frac{1}{N}\sum_{n=1}^{N}l(\theta_{n},\mathbf{N}(y_{n},\omega)).$ (5)
@@
8:use numerical optimization to solve $\omega^{*}=\underset{\omega\in\Omega}{\operatorname{argmin}}\frac{1}{N}\sum_{n=1}^{N}l(\theta_{n},\mathbf{N}(y_{n},\omega))$
@@ 3.1 Dimension-reduced Reconstruction Map (RM-DR) Estimator
$\widehat{\theta}_{RMDR}(s)=\mathbf{N}(s,\omega^{*}),\;\text{ where }\omega^{*}=\underset{\omega\in\Omega}{\operatorname{argmin}}\frac{1}{N}\sum_{n=1}^{N}l(\theta_{n},\mathbf{N}(s_{n},\omega)).$ (6)
@@
9:use numerical optimization to solve $\omega^{*}=\underset{\omega\in\Omega}{\operatorname{argmin}}\frac{1}{N}\sum_{n=1}^{N}l(\theta_{n},\mathbf{N}(s_{n},\omega))$
@@ 3.2 Connection with Bayes Estimation
In this section, we establish a connection between the RM-DR estimator and the Bayes estimator. As discussed earlier, the dimension-reduced data, $s\in\mathbb{R}^{K}$, is used as input in the reconstruction map to estimate $\theta\in\Theta\subseteq\mathbb{R}^{d}$. For a given NN architecture, denote the set of vector-valued NN functions as $\mathcal{A}=\{\mathbf{N}(\cdot,\omega)\mid\omega\in\Omega\}$, where $\mathbf{N}(\cdot,\omega):\mathbb{R}^{K}\rightarrow\mathbb{R}^{d}$. To define a Bayes estimator, we require a prior distribution $\pi(\cdot)$ on the parameters. For an estimator $g:\mathbb{R}^{K}\rightarrow\mathbb{R}^{d}$, Bayes risk is defined as $r_{s}(\pi,g):=E^{(s,\theta)\sim p_{\pi}(s,\theta)}(l(\theta,g(s)))$, where the expectation is taken over the joint distribution with density $p_{\pi}(s,\theta)=p(s\mid\theta)\pi(\theta)$. An estimator $\widehat{\theta}_{B}$ is a Bayes estimator if it minimizes the Bayes risk among all estimators:
@@
The result is stated below and the proof provided in the supplementary material.
@@
Suppose that there exists a Bayes estimator $\widehat{\theta}_{B}$ within $\mathcal{A}$, and the design density function $d(\theta)$ and prior density function $\pi(\theta)$ agree except on a set of Lebesgue measure zero. Then the RM-DR estimator converges to a Bayes estimator almost surely when the number of synthetic validation data-parameter pairs $N\rightarrow\infty$.
A desirable property of Bayes estimators is that with respect to proper priors, they are virtually always admissible (Berger, 1985), meaning that there is no other estimator as a function of $s$ that has smaller risk for every $\theta$. For a finite training sample size, the RM-DR estimator minimizes the empirical Bayes risk over a set of functions specified by the neural network architecture. The design distribution is analogous to the prior distribution, enabling the user to incorporate prior knowledge about model parameters. The theorem shows that if the model space of the neural network is rich enough, the RM-DR estimator is equivalent to a Bayes estimator when the training sample size is sufficiently large.
@@ 3.3 Understanding Dimension Reduction
To better understand the effect of dimension reduction, for any estimator $\widehat{\theta}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$, we denote the Bayes risk $r(\pi,\widehat{\theta}):=E^{(y,\theta)}(l(\theta,\widehat{\theta}(y)))$ and $\widehat{\theta}_{O}=\underset{\widehat{\theta}}{\operatorname{argmin}}\;r(\pi,\widehat{\theta})$. And we denote $\widehat{\theta}_{RMDR}^{S}(y)=\widehat{\theta}_{RMDR}(S(y))$, $\widehat{\theta}_{B}^{S}(y)=\widehat{\theta}_{B}(S(y))$ and $g^{S}(y)=g(S(y))$. Based on (7), equivalently we can write $\widehat{\theta}_{B}^{S}=\underset{g^{S}}{\operatorname{argmin}}\;r(\pi,g^{S})$. For simplicity and without loss of generality, we assume $\widehat{\theta}_{O}$ and $\widehat{\theta}_{B}$ in (7) are both unique, and $d(\theta)$ and $\pi(\theta)$ agree except on a measure of zero. Ideally, we want our estimator to be as close as possible to $\widehat{\theta}_{O}$ that minimizes the Bayes risk among all estimators as functions of full data. And it is clear that the RM estimator converges to $\widehat{\theta}_{O}$ if the number of synthetic training data-parameter pairs $N$ goes to infinity and the neural network is sufficiently expressive. However, in the finite sample case, the approximation error between $\widehat{\theta}_{RM}$ and $\widehat{\theta}_{O}$ is positive and may be large if the data dimension is high due to the inherent difficulty of estimating a function with a large input space. On the other hand, the RM-DR estimator will converge to $\widehat{\theta}_{B}^{S}$, which minimizes the Bayes risk among all estimators based on $S(\cdot)$. Therefore in the finite sample case, the discrepancy between RM-DR estimator $\widehat{\theta}_{RMDR}^{S}$ and the Bayes estimator $\widehat{\theta}_{O}$ is composed of two types of error, one is the systematic error between $\widehat{\theta}_{B}^{S}$ and $\widehat{\theta}_{O}$, and the other is the approximation error between $\widehat{\theta}_{RMDR}^{S}$ and $\widehat{\theta}_{B}^{S}$. Compared with RM estimation, although RM-DR estimation has this systematic error, the approximation error is reduced due to a lower-dimensional input. So RM-DR is able to produce a lower aggregate error, and the overall effect becomes more pronounced as data dimension increases. In terms of degree of dimension reduction, there is a trade-off between the two types of error. Generally speaking, when the dimension $K$ of the summary statistics decreases, the systematic error of the RM-DR estimator will increase due to loss of information, but the approximation error will decrease as it becomes easier to estimate a function with smaller input space. Theoretically, an estimator based on a low-dimensional sufficient statistic would be optimal since it would incur no systematic error and produce a lower approximation error that under the use of the full data. Unfortunately, finding a low-dimensional sufficient statistic in the likelihood-free setting is typically infeasible. Therefore, in practice, the choice of summaries is usually problem-specific and requires domain knowledge. In general, it is desirable for these summaries to reflect important features of data, and also depend on the unknown model parameters. For example, one may consider marginal distribution statistics such as sample moments, quantiles and order statistics. For time/spatially indexed data, we can consider descriptive features like the number of peaks or valleys, smoothness, shape of curves, frequency, amplitude, counts, etc. Summaries of temporal or spatial dependence like auto-covariance may also be useful.

### v2

@@ 1 Introduction
Our approach shares a foundation with recent literature on what we shall call reconstruction map (RM) estimation. Rudi et al. (2022) consider parameter estimation for the FitzHugh–Nagumo model by learning the mapping from the sample space to parameter space using a deep NN trained on a large number of synthetic datasets generated from the model. Similarly, Gerber and Nychka (2021) demonstrate that neural networks can be used to learn a mapping from moderate-sized spatial fields to Gaussian process covariance parameters, offering a fast alternative to computationally demanding maximum likelihood estimation. Lenzi et al. (2023) further expand its application to intractable models with an example of parameter estimation for max-stable processes, showing its flexibility in handling highly non-Gaussian and spatially dependent data. Sainsbury-Dale et al. (2024) complement this line of work by framing the reconstruction map approach as a direct approximation of the Bayes estimator—referred to as neural Bayes estimators, and by extending it to settings with independent replicates using permutation-invariant neural networks. The effectiveness of this approach is further demonstrated through additional simulation studies, including applications to a spatial conditional extremes model. Crucially, their approach suffers from the curse of dimensionality, i.e., its estimation performance degrades quickly as the data size grows. In an application for econometric models, Creel (2017) proposes to use informative statistics as input to train a neural network, the output of which can be used directly as an estimator, or as an input to subsequent classical or Bayesian inference estimation. In a similar spirit, Rai et al. (2024) propose using a set of extreme quantiles as approximate sufficient statistics as input to a neural network for parameter estimation in the generalized extreme value (GEV) distribution. More literature on neural network-based methods for inference can be found in the recent review by Zammit-Mangion et al. (2025), which provides a comprehensive overview of methodological developments in this area. Our work builds on Creel’s approach by establishing a systematic simulation-based Reconstruction Map-dimension Reduction (RM-DR) estimation method which overcomes the fundamental problem of degraded estimation performance with data dimension. We show that under certain assumptions, the resulting estimator is asymptotically equivalent to a Bayes estimator. Through multiple numerical experiments, we show that dimension reduction is essential for estimation from large datasets. We further propose a combined parameter estimation approach that utilizes the RM-DR as a starting point in a local optimization algorithm when the likelihood is available, providing an alternative to computationally costly global optimization methods.
@@ 2.1 Neural Network Models
The key building block of the NN is the neuron as shown in the left panel of Fig. 1, which consists of $n^{in}$ input nodes/variables with the $i$th variable denoted as $x_{i}$ and associated weight parameter $w_{i}$, a bias parameter $b$, an activation function $f:\mathbb{R}\rightarrow\mathbb{R}$, and the neuron output $x^{out}$. Let $x^{in}$ and $w$ be vector representations of input variables and weights, respectively. The output $x^{out}$ is calculated as evaluating the activation function at the weighted sum of input variables,
@@
More complex models can be constructed using multi-layer NNs, where there is a sequence of hidden layers between input and output layer as shown in the right panel of Fig. 1. For an L-layer NN, denote $A_{l}$ as the matrix of all parameters in the $l$th layer, and $f_{l}$ as the activation function in $l$th layer, for $l=1,\ldots,L$. Let $\omega:=(A_{1},\ldots,A_{L})$ be parameters of the NN represented as a sequence of matrices, and $\mathbf{N}(\cdot,\omega):\mathbb{R}^{n^{in}}\rightarrow\mathbb{R}^{n^{out}}$ be the vector-valued function representation of the NN,
@@ 2.2 Reconstruction Map (RM) Estimation
$\widehat{\theta}_{RM}(y)=\mathbf{N}(y,\omega^{*}),\;\text{ where }\omega^{*}\in\underset{\omega\in\Omega}{\operatorname{argmin}}\frac{1}{N}\sum_{n=1}^{N}l(\theta_{n},\mathbf{N}(y_{n},\omega)).$ (5)
@@
8:use numerical optimization to solve $\omega^{*}\in\underset{\omega\in\Omega}{\operatorname{argmin}}\frac{1}{N}\sum_{n=1}^{N}l(\theta_{n},\mathbf{N}(y_{n},\omega))$
@@ 3.1 Dimension-reduced Reconstruction Map (RM-DR) Estimator
An important practical advantage of the proposed RM-DR estimation approach is its computational efficiency relative to likelihood-free methods such as ABC and SLE. While RM-DR requires an upfront computational cost during training (involving simulation of synthetic datasets and neural network optimization), this cost is amortized across future inferences. Once the reconstruction map is learned, estimation from a new dataset under the same generative model becomes highly efficient, requiring only a forward evaluation of the pre-trained neural network and dimension reduction. Importantly, no further simulations from the model are required during inference, making the per-dataset inference cost effectively $O(1)$ with respect to the number of model simulations. In contrast, ABC requires simulating synthetic datasets until a sufficient number of accepted samples are obtained that match the observed data within a specified tolerance, leading to a computational complexity of $O(N_{\text{post}})$, where $N_{\text{post}}$ is the desired number of posterior samples. For SLE, each likelihood evaluation requires simulating $N_{s}$ synthetic datasets to estimate the synthetic likelihood, and the optimization typically requires $N_{\text{iter}}$ iterations, resulting in a computational complexity of $O(N_{s}\cdot N_{\text{iter}})$ per dataset. Consequently, RM-DR achieves estimation at a substantially lower computational cost than ABC or SLE when inference is performed on multiple or many datasets under the same model, making it well-suited for large-scale likelihood-free inference tasks where repeated parameter estimation is required.
@@
$\widehat{\theta}_{RMDR}(s)=\mathbf{N}(s,\omega^{*}),\;\text{ where }\omega^{*}\in\,\underset{\omega\in\Omega}{\operatorname{argmin}}\,\frac{1}{N}\sum_{n=1}^{N}l(\theta_{n},\mathbf{N}(s_{n},\omega)).$ (6)
@@
9:use numerical optimization to solve $\omega^{*}\in\underset{\omega\in\Omega}{\operatorname{argmin}}\frac{1}{N}\sum_{n=1}^{N}l(\theta_{n},\mathbf{N}(s_{n},\omega))$
@@ 3.2 Connection with Bayes Estimation
In this section, we establish a connection between the RM-DR estimator and the Bayes estimator. As discussed earlier, the dimension-reduced data, $s\in\mathcal{S}\subseteq\mathbb{R}^{K}$, is used as input in the reconstruction map to estimate $\theta\in\Theta\subseteq\mathbb{R}^{d}$, where $\mathcal{S}$ denotes the support of $s$. For a given NN architecture, denote the set of vector-valued NN functions as $\mathcal{A}=\{\mathbf{N}(\cdot,\omega)\mid\omega\in\Omega\}$, where $\mathbf{N}(\cdot,\omega):\mathbb{R}^{K}\rightarrow\mathbb{R}^{d}$. To define a Bayes estimator, we require a prior distribution $\pi(\cdot)$ on the parameters. For an estimator $g:\mathbb{R}^{K}\rightarrow\mathbb{R}^{d}$, Bayes risk is defined as
$r_{s}(\pi,g):=\mathbb{E}_{(s,\theta)\sim p_{\pi}(s,\theta)}\bigl{[}l(\theta,g(s))\bigr{]},$ (7)
where the expectation is taken over the joint distribution with density $p_{\pi}(s,\theta)=p(s\mid\theta)\pi(\theta)$. An estimator $\widehat{\theta}_{B}$ is a Bayes estimator if it minimizes the Bayes risk among all estimators:
@@
Let $Q_{n}(\omega)=\frac{1}{n}\sum_{i=1}^{n}l(\theta_{i},\mathbf{N}(s_{i},\omega))$ be the training loss function, where $(\theta_{i},s_{i})_{i=1}^{n}$ are i.i.d. synthetic training data-parameter pairs from the joint distribution with density $p_{d}(s,\theta)=p(s\mid\theta)d(\theta)$. Since $(\theta_{i},s_{i})$ are random, $Q_{n}(\omega)$ is a random function of $\omega$, with randomness induced by the synthetic training data. Define the expected training loss function as $Q_{0}(\omega)=\mathbb{E}_{(s,\theta)\sim p_{d}(s,\theta)}\bigl{[}Q_{n}(\omega)\bigr{]}=\mathbb{E}_{(s,\theta)\sim p_{d}(s,\theta)}\bigl{[}l(\theta,\mathbf{N}(s,\omega))\bigr{]}$. In this section, we define the RM-DR estimator as
$\widehat{\theta}_{n}(\cdot)=\mathbf{N}(\cdot,\widehat{\omega}_{n}),\quad\text{where}\quad\widehat{\omega}_{n}\in\underset{\omega\in\Omega}{\operatorname{argmin}}\;Q_{n}(\omega).$
The following theorems formalize the connection between RM-DR and Bayes estimation. Detailed proofs are provided in the supplement.
@@
Assume that:
  1. 1.
    The space $\Omega$ of parameters defining the neural network is compact;
  2. 2.
    The neural network function $\mathbf{N}(s,\omega)$ is continuous in $\omega$ for any fixed $s\in\mathcal{S}$;
  3. 3.
    The expected training loss function $Q_{0}(\omega)<\infty$ for any $\omega\in\Omega$ and has a set of minimizers $\Omega_{0}=\underset{\omega\in\Omega}{\operatorname{argmin}}\,Q_{0}(\omega)$ such that for any $\omega_{a},\omega_{b}\in\Omega_{0}$, $\mathbf{N}(\cdot,\omega_{a})=\mathbf{N}(\cdot,\omega_{b}).$ That is, the induced NN function at the minimizers is unique, denoted as $\mathbf{N}_{0}(\cdot)$;
  4. 4.
[…]
Then, the RM-DR estimator $\widehat{\theta}_{n}(\cdot)$ converges pointwise in probability to the function $\mathbf{N}_{0}(\cdot)$ as the number of synthetic training data-parameter pairs $n\to\infty$. That is, for each fixed $s\in\mathcal{S}$:
[…]
Moreover, if we additionally assume that:
  1. 5.
    The support of summary statistics $\mathcal{S}$ is compact.
  2. 6.
    The neural network function $\mathbf{N}(s,\omega)$ is jointly continuous in $(s,\omega)\in\mathcal{S}\times\Omega$,
then the RM-DR estimator $\widehat{\theta}_{n}(\cdot)$ converges uniformly in probability to the function $\mathbf{N}_{0}(\cdot)$ over $\mathcal{S}$ as $n\to\infty$. That is,
[…]
Theorem 2.
Suppose there exists a Bayes estimator $\widehat{\theta}_{B}(\cdot)$ within $\mathcal{A}=\{\mathbf{N}(\cdot,\omega):\omega\in\Omega\}$, and that the design density $d(\theta)$ and prior density $\pi(\theta)$ agree except on a set of Lebesgue measure zero. Further, suppose that Assumptions 1–4 in Theorem 1 hold. Then the RM-DR estimator $\widehat{\theta}_{n}(\cdot)$ converges pointwise in probability to the Bayes estimator $\widehat{\theta}_{B}(\cdot)$ as the number of synthetic training data-parameter pairs $n\to\infty$. That is, for each fixed $s\in\mathcal{S}$:
[…]
Moreover, if Assumptions 5-6 in Theorem 1 are also satisfied, then the RM-DR estimator $\widehat{\theta}_{n}(\cdot)$ converges uniformly in probability to the Bayes estimator $\widehat{\theta}_{B}(\cdot)$ as $n\to\infty$. That is,
[…]
Clearly,
$Q_{0}(\omega)=\mathbb{E}_{(s,\theta)\sim p_{d}(s,\theta)}\bigl{[}l(\theta,\mathbf{N}(s,\omega))\bigr{]}=r_{s}(d,\mathbf{N}(\cdot,\omega))$
can be viewed as the Bayes risk for the estimator $\mathbf{N}(\cdot,\omega)$ with the design distribution $d(\cdot)$ playing the role of the prior. One implication of Theorem 1 is that under mild assumptions, the RM-DR estimator will converge in probability to an estimator that minimizes the Bayes risk over the class of functions specified by the neural network architecture. For a finite training sample size, the RM-DR estimator minimizes the empirical Bayes risk over the specified neural network class, with the design distribution serving as an analogue to the prior, thereby allowing the incorporation of prior knowledge about the parameter distribution.
A desirable property of Bayes estimators is that with respect to proper priors, they are virtually always admissible (Berger, 1985), meaning that there is no other estimator, as a function of $s$, that achieves a strictly smaller risk for every $\theta$. Theorem 2 then shows that if the neural network class is sufficiently rich, under mild conditions, the RM-DR estimator becomes equivalent to the Bayes estimator as the training sample size grows large. This connection provides a theoretical justification for using RM-DR in practice, ensuring that as the neural network class becomes sufficiently expressive and the training sample size grows, RM-DR yields Bayes-optimal decisions under mild assumptions.
@@ 3.3 Understanding Dimension Reduction
To better understand the effect of dimension reduction, for any estimator $\widehat{\theta}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$, we denote the Bayes risk $r(\pi,\widehat{\theta}):=\mathbb{E}_{(y,\theta)}\bigl{[}l(\theta,\widehat{\theta}(y))\bigr{]}$ and $\widehat{\theta}_{O}=\underset{\widehat{\theta}}{\operatorname{argmin}}\;r(\pi,\widehat{\theta})$. And we denote $\widehat{\theta}_{RMDR}^{S}(y)=\widehat{\theta}_{RMDR}(S(y))$, $\widehat{\theta}_{B}^{S}(y)=\widehat{\theta}_{B}(S(y))$ and $g^{S}(y)=g(S(y))$. Based on (8), equivalently we can write $\widehat{\theta}_{B}^{S}=\underset{g^{S}}{\operatorname{argmin}}\;r(\pi,g^{S})$. For simplicity and without loss of generality, we assume $\widehat{\theta}_{O}$ and $\widehat{\theta}_{B}$ in (8) are both unique, and $d(\theta)$ and $\pi(\theta)$ agree except on a measure of zero. Ideally, we want our estimator to be as close as possible to $\widehat{\theta}_{O}$ that minimizes the Bayes risk among all estimators as functions of full data. And it is clear that the RM estimator converges to $\widehat{\theta}_{O}$ if the number of synthetic training data-parameter pairs $N$ goes to infinity and the neural network is sufficiently expressive. However, in the finite sample case, the approximation error between $\widehat{\theta}_{RM}$ and $\widehat{\theta}_{O}$ is positive and may be large if the data dimension is high due to the inherent difficulty of estimating a function with a large input space. On the other hand, the RM-DR estimator will converge to $\widehat{\theta}_{B}^{S}$, which minimizes the Bayes risk among all estimators based on $S(\cdot)$. Therefore in the finite sample case, the discrepancy between RM-DR estimator $\widehat{\theta}_{RMDR}^{S}$ and the Bayes estimator $\widehat{\theta}_{O}$ is composed of two types of error, one is the systematic error between $\widehat{\theta}_{B}^{S}$ and $\widehat{\theta}_{O}$, and the other is the approximation error between $\widehat{\theta}_{RMDR}^{S}$ and $\widehat{\theta}_{B}^{S}$. Compared with RM estimation, although RM-DR estimation has this systematic error, the approximation error is reduced due to a lower-dimensional input. So RM-DR is able to produce a lower aggregate error, and the overall effect becomes more pronounced as data dimension increases. In terms of degree of dimension reduction, there is a trade-off between the two types of error. Generally speaking, when the dimension $K$ of the summary statistics decreases, the systematic error of the RM-DR estimator will increase due to loss of information, but the approximation error will decrease as it becomes easier to estimate a function with smaller input space. Theoretically, an estimator based on a low-dimensional sufficient statistic would be optimal since it would incur no systematic error and produce a lower approximation error that under the use of the full data. Unfortunately, finding a low-dimensional sufficient statistic in the likelihood-free setting is typically infeasible. Therefore, in practice, the choice of summaries is usually problem-specific and requires domain knowledge. In general, it is desirable for these summaries to reflect important features of data, and also depend on the unknown model parameters. For example, one may consider marginal distribution statistics such as sample moments, quantiles and order statistics. For time/spatially indexed data, we can consider descriptive features like the number of peaks or valleys, smoothness, shape of curves, frequency, amplitude, counts, etc. Summaries of temporal or spatial dependence like auto-covariance may also be useful.